OpenStudy (anonymous):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f x x a b g x c x d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

anyone please

OpenStudy (anonymous):

can you draw what the functions look like?

OpenStudy (anonymous):

f(x)=x+a/c g(x)=cx+d

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

x+a/b my bad

OpenStudy (anonymous):

For f(x) which one of these is it?